OpenStudy (anonymous):

Choose the correct simplification of the expression (5a^7b)^2

OpenStudy (anonymous):

25a^14b^2 25a^9b^2 25a^14b 10a^14b^2

OpenStudy (anonymous):

the third one

OpenStudy (ajprincess):

use \(\large(a^b)^c=a^{bc}\) \(\large(5a^7b)^2=5^2*a^{7*2}*b^2\) Can u do nw?

jhonyy9 (jhonyy9):

so what is your opinion ?

jhonyy9 (jhonyy9):

do you like to cooperate now ?

jhonyy9 (jhonyy9):

@ajprincess have wrote it right sure to you

jhonyy9 (jhonyy9):

can you do it

jhonyy9 (jhonyy9):

???

jhonyy9 (jhonyy9):

you need make just the calcules

OpenStudy (anonymous):

@sommersquid you mean \[(5a^{7b})^{2}\] or \[(5a^{7}b)^{2}\] ?

OpenStudy (anonymous):

Second one ^

OpenStudy (ajprincess):

hmm no. plzz try again

OpenStudy (anonymous):

Well for this you can just represent the square as two things being multiplied

OpenStudy (ajprincess):